use two unit multipliers to convert 1000 square centimeters to square meters.
Found 2 solutions by stanbon, Mathtut:
Answer by stanbon(75887)

use two unit multipliers to convert 1000 square centimeters to square meters.
-----------------------
1000 cm^2(1 meter/100 cm)(1 meter/100 cm) = 1000/(100)^2 square meters
= 1000/10000 sq. meters = 1/10 sq. meters
===========================================
Cheers,
Stan H.

Answer by Mathtut(3670)